OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to given line (0,4);9x+8y=5

OpenStudy (anonymous):

9x+8y=5 y=(-9/8)x+5/8 make the line y=(8/9)X+B Through 0,4 So 4=B So Y=(8/9)x+4

OpenStudy (anonymous):

ur my hero thnks